Two dozen eggs and a loaf of bread coast K4.82.Half a dozen eggs and two loaves of bread cost K2.29.Find the cost of a loaf of bread.
Answer by stanbon(75887)

Two dozen eggs and a loaf of bread coast K4.82.Half a dozen eggs and two loaves of bread cost K2.29.Find the cost of a loaf of bread.
---------
Equations:
2e + b = 4.82
(1/2)e+2b = 2.29
---------
Modify and solve for "b"::
2e + b = 4.82
2e + 8b = 4*2.29
----
Subtract and solve for "b"::
7b = 4.34
b = K0.62 per loaf
-----------
Cheers,
Stan H.
---------------